Accelerator Update for May 28, 2008 - May 30, 2008
Wednesday May 28
The day shift began with Operations monitoring the Recycler stash (139.1E10) and the Antiproton Source (Pbar) stack (14.0E10), with Pbar not stacking due to the F23 Yard Breaker work, with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) waiting for beam, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:30 AM, a Booster RF station (BRF4) tripped off and would not reset. Experts investigated.
FESS reported at 9:02 AM that the F23 Yard Breaker work was completed.
Pbar resumed stacking at 9:37 AM.
At 10:28 AM, LRF3 tripped off. It reset.
A Booster expert reported around 11:40 AM that he was working on BRF4 and BRF13. Both stations were down with a bias supply problem.
At 11:52 AM, the NuMI beam permit tripped off due to an incomplete RF gap. Everything reset.
Operations put the TeV through a dry squeeze at 12:51 PM.
Operations put the TeV into shot setup at 1:15 PM.
Operations reported at 1:20 PM that all the Booster RF stations were working. BRF19 was returned to standby mode.
Operations established TeV store 6174 at 2:57 PM with an initial luminosity of 191.3E30.
Operations reported at 3:12 PM that a Recycler expert was going to turn off the damper system for 30 minutes to conduct maintenance.
Operations reported at 3:14 PM that a MI expert had completed his intensity variant measurements.
Operations reported at 4:04 PM that the Recycler had stashed antiprotons four times during the day shift.
At 4:39 PM, Operations reported a smaller set of transfers of antiprotons than normal to the Recycler. The Recycler VSA had a frequency width problem.
At 5:21 PM, the BNB beam permit tripped off due to the current drifting low on a power supply (E:HV860). Everything reset.
At 5:49 PM, Operations reported a Pbar readback problem with their BPMs during the last stash. Operators reset the BPMs, re-armed them, and then reloaded the trigger events.
At 8:08 PM, Operations received a Meson “MS2 LCW Pressure” alarm. Later, operators investigated and found a leak on the downstream end of quadrupole magnet (MT4Q3). They valved out the magnet.
At 9:50 PM, a Booster vacuum pump (IP61) tripped off. It reset.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Thursday May 29
The midnight shift began with Operations monitoring store 6174 (68.98E30) and the stash (228.8E10), with Pbar stacking (37.5E10), and with NuMI and BNB taking beam.
Operations terminated store 6174 at 2:24 AM.
Operations put the TeV into shot setup at 2:35 AM.
At 3:49 AM, BRF10 tripped off. It reset.
Operations established store 6175 at 4:19 AM with an initial luminosity of 222.26E30.
Operations reported at 7:55 AM that the Recycler had stashed five times during the midnight shift.
MI and Booster experts began an 8 GeV optics study of the MI-8 transfer line at 9:48 AM.
At 9:50 AM, Operations reported a Pbar sequencer error.
An instrumentation expert reported at 10:16 AM that he was rebooting the MI DCCT FE.
At 10:36 AM, Operations reported a chilled water leak in the Cross Gallery’s Computer Room cooling system. Experts will install two 10 ton spot coolers, turn the cooling system off, and repair the leak.
At 11:09 AM, a Linac quadrupole power supply (LPSQ103) tripped off for the first of around five times in about 30 minutes. Linac experts had the power supply repaired by 11:40 AM.
A MI expert began a coalescing study at 1:49 PM.
At 2:05 PM, the 400MeV area beam permit tripped off due to bad beam positions. It reset.
FESS personnel had the Cross Gallery’s Computer Room chilled water leak repaired by 2:32 PM.
At 2:51 PM, Operations reported that the second stash from Pbar to the Recycler never made it out of MI. The MI-30 kicker fired late.
Operations reported at 3:56 PM that the Recycler had stashed five times during the day shift.
The MI expert completed her coalescing study at 4:33 PM.
At 6:55 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to also trip off. Everything reset.
Operations terminated store 6175 at 7:09 PM.
Operations put the TeV into shot setup at 7:32 PM.
Operations established store 6176 at 9:01 PM with an initial luminosity of 240.15E30.
At 11:07 PM, LRF1 tripped off. It reset.
At 11:19 PM, Operations reported that a Linac quadrupole’s power supply (L:QPS101) current was drifting. Operators swapped cards and provided extra cooling. Operations contacted a Linac expert.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Friday May 30
The midnight shift began with Operations monitoring store 6176 (131.35E30) and the stash (85.6E10), with Pbar stacking (35.5E10), and with NuMI and BNB taking beam.
Around 1 AM, Operations reported a 5% drop in luminosity. The tune went erratic for a few seconds and then returned to normal. Operators reported that a vertical tune change had occurred about 1.5 minutes before this happened.
At 2:39 AM, Operations reported that there had been a number of MI aborts due to L:QPS101’s current drift. Operations reduced the Linac intensity while operators investigated.
Operations reported that operators had brought the Linac intensity back to normal by 3:09 AM.
Operations reported at 7:55 AM that the Recycler had stashed six times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 05/30/2008 email Fermilab|
|Security, Privacy, Legal|